OpenStudy (anonymous):

Solve using the Quadratic Formula. 2x2 + 3x - 6 = 0

OpenStudy (anonymous):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \]

OpenStudy (anonymous):

in this case can you tell me what your a b c are?

OpenStudy (anonymous):

2 is my a 3x is my b and 6 is my c

OpenStudy (anonymous):

umm coefficients. We don't want any x's anyways you got a right b you just need the 3 c you forgot the - sign

OpenStudy (anonymous):

a=1 b=3 c =-6 You plug that in to the first equation but you have to do it twice. What do I mean by doing it twice? Well do you see the \[\pm\] This signs means plus or minus. So you have to solve it with + and then you have to solve it again for -

OpenStudy (anonymous):

\[x=1/4(-3-\sqrt{57})\] \[x=1/4(\sqrt{57}-3)\]

OpenStudy (anonymous):

@pariisz241 Thank you but it's pretty straight forward from where I left off unless the person doesn't know how to plug in values or how to solve for x. Good job.